If a number is added to its square the result is 56. find the number.

thank you

Answer by fractalier(6550)

Let the number be x. So we have
x + x^2 = 56
x^2 + x - 56 = 0
(x + 8)(x - 7) = 0
x = -8 or x = 7
Two answers and both work.